# Data Distribution-based Curriculum Learning

### v1

@@ Abstract
The order of training samples can have a significant impact on the performance of a classifier. Curriculum learning is a method of ordering training samples from easy to hard. This paper proposes the novel idea of a curriculum learning approach called Data Distribution-based Curriculum Learning (DDCL). DDCL uses the data distribution of a dataset to build a curriculum based on the order of samples. Two types of scoring methods known as DDCL (Density) and DDCL (Point) are used to score training samples thus determining their training order. DDCL (Density) uses the sample density to assign scores while DDCL (Point) utilises the Euclidean distance for scoring. We evaluate the proposed DDCL approach by conducting experiments on multiple datasets using a neural network, support vector machine and random forest classifier. Evaluation results show that the application of DDCL improves the average classification accuracy for all datasets compared to standard evaluation without any curriculum. Moreover, analysis of the error losses for a single training epoch reveals that convergence is faster when using DDCL over the no curriculum method.
@@
data distribution, curriculum learning, scoring method, classification, neural network, support vector machine, random forest.
@@ I Introduction
To reduce the negative impact of these factors, the training algorithms applied to datasets are often designed to guide the learning model towards optimum performance [8, 9, 10]. Gradient descent is the most common method of optimising a neural network due to its fast convergence towards the minimum error [11]. Variants of gradient descent such as batch gradient descent, stochastic gradient descent (SGD) and mini batch gradient descent exist for use with specific problem scenarios. These problems may favour computation speed, data size or a balance of speed and size as frequently seen in practice [12]. Support vector machines are optimised by selecting a value for the C hyperparameter and the method of optimising its kernel varies according to the kernel used. The Radial Basis Function (RBF) kernel is commonly used in a SVM and it is optimised by selecting the gamma ($\gamma$) hyperparameter. Both of these hyperparameters can be found using several search algorithms including trees of Parzen estimators, particle swarm optimization and Bayesian optimisation. Recent findings have concluded that the first two algorithms provide better hyperparameter values with a lower execution time compared to Bayesian optimisation which has a very high computational cost [13]. In a random forest classifier, the number of decision trees (estimators) and the number of features required to split a tree are hyperparameters that are optimised [14].
@@
Furthermore, the data distribution plays a major role in obtaining quality predictions from classifiers. When using the same data, a particular data distribution of training data can provide better results compared to another data distribution. Research has shown that the choice of distribution can cause an overall low bias or high bias for the same training data and also lead to poor performance without being apparent [15, 16]. This bias can have a significant impact on the model’s performance which can be critical in medical applications [15, 17].
Once the distribution of data is known, a decision can be made on how data samples are selected from the dataset for optimal results. This decision may include using a specific criteria for selecting samples in a particular order. The process of selectively choosing samples from a dataset as well as determining their order is known as creating a curriculum.
@@
Using a curriculum for improving performance of a machine learning model is known as curriculum learning and it is typically applied on the order of training data samples. Curriculum learning enhances training of a model by starting with simple concepts and then gradually introducing difficult concepts as training progresses [18]. It is inspired by how humans are taught through curriculums in an education system by receiving basic education during childhood and then moving on to advanced education in adulthood [19, 20].
@@
Curriculum learning has been applied to various problems achieving excellent results in the areas of image classification [21, 22], face recognition [23], visual attribute classification [24] and imbalanced data classification [25]. It has also been successfully used to improve performance of existing systems that use machine learning including unsupervised domain adaptation [26], transfer learning [27, 28] and reinforcement learning [29, 30].
@@
In this paper, we propose a data distribution based curriculum learning approach for classification tasks. The proposed approach first determines the distribution of data and then information from the distribution is used to build a curriculum based on the order of samples. The approach is then evaluated on multiple datasets with classifiers based on three different types of widely used learning methods: neural networks, SVM and random forest classifier. The remainder of the paper is organised as follows. Section II provides a background on curriculum learning and related learning approaches as well as applications of curriculum learning to various problems. The proposed curriculum learning approach is discussed in Section III and Section IV outlines its experimental evaluation. Section V examines the experiment results while Section VI provides concluding remarks.
@@ II Background
Kumar et al. [31] built upon the concepts introduced by Bengio et al. by proposing self-paced learning (SPL), a training strategy that also suggested presenting training samples ordered from simple to complex. The main difference between SPL and curriculum learning was the process of defining the difficulty of a sample. In SPL, the current learning progress of a model was considered for selecting the next sample whereas curriculum learning relied on a pre-determined fixed curriculum. The authors conducted experiments that showed their proposed method outperforming the state of the art method.
Years later, Jiang et al. proposed a framework called self-paced curriculum learning (SPCL) that sought to unify the ideas of curriculum learning and self-paced learning [32]. They reasoned that both types of learning have drawbacks; The curriculum in curriculum learning is heavily reliant on the quality of fixed pre-determined prior knowledge and ignores feedback about the learner. In self-paced learning, the curriculum is dynamically determined to adjust to the learning pace of the learner but is prone to over-fitting due to being unable to handle prior knowledge. According to the authors, SPCL addresses these drawbacks by introducing a flexible way of including prior knowledge while also dynamically adjusting the curriculum based on feedback from the learner.
Hacohen et al. [28] expanded the definition of curriculum learning by describing it as consisting of two tasks with specific functions termed as the scoring function and the pacing function. The scoring function ranks the difficulty of the samples in a curriculum while the pacing function determines how often newer samples are presented to a model during training.
@@
Overall, we categorise the reviewed curriculum learning literature into three different groups: the applications of standard curriculum learning [24, 30, 17], variations of learning styles [27, 28, 23, 22] and research that makes use of the data density [21, 26, 25].
@@ II-A Standard curriculum learning
Sarafianos et al. [24] applied curriculum learning to visual attribute classification by introducing a method that combined curriculum learning with multi-task learning. Their method performed end-to-end learning by providing a ConvNet with a complete image of a human without additional data to aid in classification. Then in multi-task learning, the tasks were split into groups using clustering. Curriculum learning was applied to the groups starting with the highest within-group cross-correlation and moving to the less correlated ones. Experiments performed on three datasets consisting of humans standing resulted in an increase in performance by up to 10%.
Tidd et al. [30] use curriculum learning to train deep reinforcement learning policies for bipedal walking of a robot over challenging terrain. The authors created a easy to hard curriculum using a three stage framework: In the first stage, guiding forces were applied to the joints and the base of the robot to start learning on easy terrain and gradually increased to difficult terrain. At the second stage when the terrain was most difficult, the guiding forces applied to the robot were slowly decreased. During the final stage, the magnitude of external random perturbations were increased to improve the robustness of the policy. Simulation experiments conducted by the authors demonstrated that a curriculum approach was effective in learning to walk for five types of terrain.
@@
Figure 1: Data Distribution-based Curriculum Learning
@@ II-B Learning style variations
In [27], the Weinshall et al. used transfer learning to implement curriculum learning. They presented an approach where the curriculum was inferred through transfer learning from another network that was pre-trained on another task. A CNN with two architectures trained using curriculum learning was used to evaluate the proposed method on the CIFAR-100 and STL-10 datasets. Experiment results from their work concluded that with curriculum learning the convergence is faster during the beginning of training and improved generalisation is achieved when more difficult tasks are used.
Hacohen et al. applied curriculum learning to the training of deep networks [28]. They used transfer learning and bootstrapping as two different techniques for the easy to hard process which they termed the scoring function. With transfer learning, a ‘teacher’ network was trained on a large dataset and its prediction performance was used as the scoring function. In bootstrapping, a network was trained without a curriculum and its performance on the training data defined the scoring function. Then, the network was retrained from scratch using curriculum learning. For the pacing function, three approaches were used with all pacing functions having comparable performance based on experiments done by the authors. Experiments done on 6 test cases showed that using a curriculum provided high accuracy and better convergence with the amount of improvement ranging from small to large depending on the test case.
@@
Recent research has looked at improving the general curriculum learning process. Zhou et al. presented a curriculum learning method known as dynamics-optimized curriculum learning (DoCL) [22]. DoCL selected training samples at each step using weighted sampling based on the scores. The authors conducted experiments on more than 9 datasets achieving results that significantly improved the performance and efficiency compared to existing curriculum learning methods.
@@ II-C Data density based curriculum learning
Our review of literature for curriculum learning has found few papers that utilise the density of the data. These works are presented in this section and our work on DDCL adds to this field of research.
Applying curriculum learning to large data sets has shown promising results. Researchers in [21] presented a method for training deep neural networks on large-scale weakly-supervised web images. Their training strategy utilised curriculum learning to effectively handle the large amount of noisy labels and data imbalance during the training process. They proposed a curriculum created by measuring the complexity of data using the density of its data. Experiments using the proposed training strategy resulted in state-of-the-art performance on benchmark data sets that included Webvision, ImageNet, Clothing-1M and Food-101.
@@
// Variables:
// Classes: $S$
// Data size: $N$
// $N_{s}$ data set for class $s$
// $|N_{s}|$ data size (cardinality)
// $h$ bandwidth
1 foreach $s\in S$ do
2       ${O_{s}}=\sum_{n=0}^{|N_{s}|}\min_{\mu_{s}\in S}(\parallel x_{n}-\mu_{s}\parallel^{2})$
3       $s=s\cup O_{s}$
4       $E_{s}=\varnothing$
5       foreach $x\in s$ do
6             $E_{s}=E_{s}\>\cup\parallel{O_{s}-x}\parallel$
7
8       end foreach
9      $D_{s}=\frac{1}{|N_{s}|h}\sum_{i=1}^{|N_{s}|}K(\frac{y-\hat{E}_{i}}{h})$
10       for $q\in Q^{s}$ do
11             if $|q|<3$ then
12                   $q=SMOTE(q)$
13             end if
14
15       end for
16      if Scoring type normal-$\leftarrow$ “Density” then
17            go to Algorithm 2
18       else if Scoring type normal-$\leftarrow$ “Point” then
19            go to Algorithm 3
20       end if
      Result: $T\leftarrow R_{s}$
21
22 end foreach
@@ III Proposed Curriculum Learning Approach
The curriculum learning approach proposed in this paper is based on the data distribution of a dataset. This is referred to as Data Distribution-based Curriculum Learning (DDCL) and involves multiple steps as visualised through Figure 1. It begins with dividing the data into groups based on their target classes. Then, the centroids for each group is calculated (Centroid Determination) and used to compute the Euclidean distance between each sample and its centroid. The data distribution for every class is determined next and utilised to divide the data for individual classes into quantiles (Q1, Q2, Q3). After quantile division, oversampling is optionally performed on quantiles where the number samples may be highly imbalanced compared to other quantiles. Lastly, the data is scored with either density or point scoring and rearranged for use in training.
The DDCL process is further detailed in Algorithm 1. In the first step of DDCL, the training data samples are grouped according to their classes ($S$) with the number of groups being equal to the number of unique classes in the data. Next, the grouped training data is used to calculate the centroids ($O_{s}$) for each class by performing clustering on each data group. Then, the centroids are used to compute the Euclidean distance from the relevant centroid ($E_{s}$) for each data sample ($x$). The third step determines the data distribution for a class ($D_{s}$) by plotting the normalised values of the Euclidean distances. Once the distribution of the training data is known, the data samples are divided into quantiles ($Q^{s}$). Each quantile ($q$) is then examined for the number of samples and the quantiles with the lowest number of samples are oversampled using SMOTE [33] if there are sufficient samples. SMOTE creates synthetic samples using the original data and different variants of it have been developed [34] since the initial method was proposed. The initial method of SMOTE is used in our proposed approach to address the potential lack of samples in a given quantile. Otherwise, no oversampling is applied due to insufficient samples. After the samples are divided into quantiles and optionally oversampled, each data sample is scored using either density or point scoring and rearranged accordingly ($R_{s}$). Finally, the rearranged training data ($T$) is ready to be passed to one of the learning methods for training and evaluation.
(a) Density scoring
(b) Point scoring
Figure 2: DDCL scoring methods
DDCL consists of two types of scoring methods: a sample density based method and a Euclidean distance based method. These are referred to as DDCL (Density) and DDCL (Point) respectively and are detailed in the following sections.

### v2

@@ Abstract
The order of training samples can have a significant impact on a model’s performance. Curriculum learning is an approach for gradually training a model by ordering samples from ‘easy’ to ‘hard’. This paper proposes the novel idea of a curriculum learning strategy called Data Distribution-based Curriculum Learning (DDCL). DDCL uses the inherent data distribution of a dataset to build a curriculum based on the order of samples. Our proposed approach is innovative as it incorporates two distinct scoring methods known as DDCL-Density and DDCL-Point to determine the order of training samples. The DDCL-Density method assigns scores based on the density of samples favoring denser regions that can make initial learning easier. Conversely, DDCL-Point utilizes the Euclidean distance from the centroid of the dataset as a reference point to score samples providing an alternative perspective on sample difficulty. We evaluate the proposed DDCL approach by conducting experiments across various classifiers using a diverse set of small to medium-sized medical datasets. Results show that DDCL improves the classification accuracy, achieving increases ranging from 2% to 10% compared to baseline methods and other state-of-the-art techniques. Moreover, analysis of the error losses for a single training epoch reveals that DDCL not only improves accuracy but also increases the convergence rate, underlining its potential for more efficient training. The findings suggest that DDCL can specifically be of benefit to medical applications where data is often limited and indicate promising directions for future research in domains that involve limited datasets.
@@
classification, curriculum learning, data distribution, machine learning, neural network, random forest, support vector machine
[…]
=-21pt
@@ I Introduction
To reduce the negative impact of these factors, the training algorithms applied to datasets are often designed to guide the learning model towards optimum performance [8, 9, 10]. Guidance for smaller datasets is particularly important since their limited size cannot provide the sample diversity present in larger datasets [11].
@@
Gradient descent is the most common method of optimizing a neural network due to its fast convergence towards the minimum error [12]. Variants of gradient descent such as batch gradient descent, stochastic gradient descent (SGD) and mini-batch gradient descent exist for use with specific problem scenarios [13]. These problems may favor computation speed, data size or a balance of speed and size as frequently seen in practice [14]. Support vector machines are optimized by selecting a value for the C hyperparameter and the method of optimizing its kernel varies according to the kernel used. The Radial Basis Function (RBF) kernel is commonly used in a SVM and it is optimized by selecting the gamma ($\gamma$) hyperparameter. Both of these hyperparameters can be found using several search algorithms including trees of Parzen estimators, particle swarm optimization and Bayesian optimization. Recent findings have concluded that the first two algorithms provide better hyperparameter values with a lower execution time compared to Bayesian optimization which has a very high computational cost [15]. In a random forest classifier, the number of decision trees (estimators) and the number of features required to split a tree are hyperparameters that are optimized [16]. The number of estimators for a specific problem is carefully selected since it corresponds with the computation time required.
@@
Furthermore, the data distribution plays a major role in obtaining quality predictions from classifiers. When using the same data, a particular data distribution of training data can provide better results compared to another data distribution. Research has shown that the choice of distribution can cause bias in the training data due to under-representation of the minority class [17]. This bias can have a significant impact on a model’s accuracy and precision which can be critical for medical applications where training data can be limited [18, 19, 20]. Once the distribution of data is known, a decision can be made on how data samples are selected from the dataset for optimal results. This decision may include using a specific criteria for selecting samples in a particular order. The process of selectively choosing samples from a dataset as well as determining their order is known as creating a curriculum.
@@
Using a curriculum for improving performance of a machine learning model is known as curriculum learning [21] and it is typically applied on the order of training data samples. Curriculum learning enhances training of a model by starting with simple concepts and then gradually introducing difficult concepts as training progresses [21, 22]. It is inspired by how humans are taught through curriculums in an education system by receiving basic education during childhood and then moving on to advanced education in adulthood [23, 24].
Curriculum learning has been applied to various problems achieving excellent results in the areas of image classification [25, 26], face recognition [27], visual attribute classification [28] and imbalanced data classification [29]. The effectiveness of the ‘easy’ to ‘hard’ strategy, where simpler examples are presented before more complex ones, has been well-documented in literature [30, 31].
However, rival approaches offer different strategies by focusing on sample difficulty and order during training. Self-paced learning (SPL) [32], a training strategy that also suggested presenting training samples ordered from simple to complex, built upon the concepts introduced in curriculum learning by altering the process of defining the difficulty of a sample. In SPL, the current learning progress of a model was considered for selecting the next sample whereas curriculum learning relied on a pre-determined fixed curriculum.
The concepts of SPL were used years later in a framework called self-paced curriculum learning (SPCL) [33]. SPCL sought to unify the ideas of curriculum learning and SPL. The authors of the study reasoned that both types of learning have drawbacks; The curriculum in curriculum learning is heavily reliant on the quality of fixed pre-determined prior knowledge and ignores feedback about the learner. In SPL, the curriculum is dynamically determined to adjust to the learning pace of the learner but is prone to over-fitting due to being unable to handle prior knowledge. According to the authors, SPCL addressed these drawbacks by introducing a flexible way of including prior knowledge while also dynamically adjusting the curriculum based on feedback from the learner.
@@
Despite this, curriculum learning has been successfully used to improve performance of existing systems that use machine learning including unsupervised domain adaptation [34], transfer learning [35, 36] and reinforcement learning [30, 37].
@@
In this paper, we propose a data distribution based curriculum learning approach for classification tasks on small to medium-sized medical datasets. The proposed approach first determines the distribution of data and then information from the distribution is used to build a curriculum based on the order of samples. The approach is then evaluated on multiple datasets with classifiers based on three different types of widely used learning methods: neural networks, SVM and random forest classifier. The remainder of the paper is organized as follows. Section II provides a background on curriculum learning as well as applications of curriculum learning to various problems. The proposed curriculum learning approach is discussed in Section III and Section IV outlines its experimental evaluation. Section V examines the experiment results whereas Section VI provides concluding remarks.
@@ II Background
Hacohen et al. [36] expanded the definition of curriculum learning by describing it as consisting of two tasks with specific functions, termed the scoring function and the pacing function. The scoring function ranks the difficulty of the samples in a curriculum whereas the pacing function determines how often newer samples are presented to a model during training.
@@
Overall, we categorize the reviewed curriculum learning literature into three different groups: the applications of standard curriculum learning [28, 37, 19], variations of learning styles [35, 36, 27, 26] and research that makes use of the data density [25, 34, 29].
@@ II-A Standard curriculum learning
Sarafianos et al. [28] applied curriculum learning to visual attribute classification by introducing a method that combined curriculum learning with multi-task learning. Their method performed end-to-end learning by providing a convolutional neural network (CNN) with a complete image of a human without additional data to aid in classification. Then in multi-task learning, the tasks were split into groups using clustering. Curriculum learning was applied to the groups starting with the highest within-group cross-correlation and moving to the less correlated ones. Experiments performed on three datasets consisting of humans standing resulted in an increase in performance by up to 10%.
Tidd et al. [37] use curriculum learning to train deep reinforcement learning policies for bipedal walking of a robot over challenging terrain. The authors created an easy to hard curriculum using a three stage framework: In the first stage, guiding forces were applied to the joints and the base of the robot to start learning on easy terrain and gradually increased to difficult terrain. At the second stage when the terrain was most difficult, the guiding forces applied to the robot were slowly decreased. During the final stage, the magnitude of external random perturbations were increased to improve the robustness of the policy. Simulation experiments conducted by the authors demonstrated that a curriculum approach was effective in learning to walk for five types of terrain.
Figure 1: Data Distribution-based Curriculum Learning
@@ II-B Learning style variations
In [35], the Weinshall et al. used transfer learning to implement curriculum learning. They presented an approach where the curriculum was inferred through transfer learning from another network that was pre-trained on another task. A CNN with two architectures trained using curriculum learning was used to evaluate the proposed method on the CIFAR-100 [38] and STL-10 datasets [39]. Experiment results from their work concluded that with curriculum learning the convergence is faster during the beginning of training and improved generalization is achieved when more difficult tasks are used.
Hacohen et al. applied curriculum learning to the training of deep networks [36]. They used transfer learning and bootstrapping as two different techniques for the easy to hard process which they termed the scoring function. With transfer learning, a ‘teacher’ network was trained on a large dataset and its prediction performance was used as the scoring function. In bootstrapping, a network was trained without a curriculum and its performance on the training data defined the scoring function. Then, the network was retrained from scratch using curriculum learning. For the pacing function, three approaches were used with all pacing functions having comparable performance based on experiments done by the authors. Experiments done on six test cases showed that using a curriculum provided high accuracy and better convergence with the amount of improvement ranging from small to large depending on the test case.
@@
Recent research has looked at improving the general curriculum learning process. Zhou et al. [26] presented a curriculum learning method known as dynamics-optimized curriculum learning (DoCL). DoCL selected training samples at each step using weighted sampling based on the scores. The authors conducted experiments on more than nine datasets achieving results that significantly improved the performance and efficiency compared to existing curriculum learning methods.
@@ II-C Data density based curriculum learning
Our review of literature for curriculum learning has found few papers that utilize the density of the data. These works are presented in this section and our work on DDCL adds to this field of research. Specifically, DDCL addresses the key limitation of other works [25, 34] that require large datasets to utilize data density by effectively operating on small datasets. Moreover, clustering in DDCL is used to determine the class centroids rather than the density compared to other works.
Researchers in [25] presented a method for training deep neural networks on large-scale weakly-supervised web images. Their training strategy utilized curriculum learning to effectively handle the large amount of noisy labels and data imbalance during the training process. They proposed a curriculum created by measuring the complexity of data using the density of its data. Experiments using the proposed training strategy resulted in state-of-the-art performance on benchmark datasets.
@@
(a) Density scoring showing the densities of each quantile.
(b) Point scoring showing samples sorted by the $\hat{E_{s}}$ value.
Figure 2: Example of DDCL scoring methods on Haberman’s Survival data.
@@
1:  Variables:
2:  Classes: $S$
3:  Data size: $N$
4:  $N_{s}$ dataset for class $s$
5:  $|N_{s}|$ data size (cardinality)
6:  $h$ bandwidth
7:  for all $s\in S$ do
8:     ${O_{s}}\leftarrow\sum_{n=0}^{|N_{s}|}\min_{\mu_{s}\in S}(\parallel x_{n}-\mu_{s}\parallel^{2})$
9:     $s\leftarrow s\cup O_{s}$
10:     $E_{s}\leftarrow\varnothing$
11:     for all $x\in s$ do
12:        $E_{s}\leftarrow E_{s}\cup\parallel{O_{s}-x}\parallel$
13:     end for
14:     $D_{s}\leftarrow\frac{1}{|N_{s}|h}\sum_{i=1}^{|N_{s}|}K\left(\frac{y-\hat{E}_{i}}{h}\right)$Kernel Density Estimation}
15:     for $q\in Q^{s}$ do
16:        if $|q|<3$ then
17:           $q\leftarrow SMOTE(q)$
18:        end if
19:     end for
20:     if Scoring type $\leftarrow$ “Density” then
21:        go to Algorithm 2
22:     else if Scoring type $\leftarrow$ “Point” then
23:        go to Algorithm 3
24:     end if
25:
26:     return  $T\leftarrow R_{s}$
27:  end for
@@ III Proposed Curriculum Learning Approach
The curriculum learning approach proposed in this paper is based on the data distribution of a dataset. This is referred to as Data Distribution-based Curriculum Learning (DDCL) and involves multiple steps as visualized through Figure 1. It begins with dividing the data into groups based on their target classes. Then, the centroids for each group are calculated (Centroid Determination) and used to compute the Euclidean distance between each sample and its centroid. The data distribution for each class is determined next and utilized to divide the data for individual classes into quantiles (Q1, Q2, Q3). After quantile division, oversampling is optionally performed on quantiles where the number of samples may be highly imbalanced compared to other quantiles. Lastly, the data is scored with either density or point scoring and rearranged for use in training.
The DDCL process is further detailed in Algorithm 1, Algorithm 2 and Algorithm 3. In the first step of DDCL, the training data samples are grouped according to their classes ($S$) with the number of groups being equal to the number of unique classes in the data. Next, the grouped training data is used to calculate the centroids ($O_{s}$) for each class by performing clustering on each data group. Then, the centroids are used to compute the Euclidean distance from the relevant centroid ($E_{s}$) for each data sample ($x$). The third step determines the data distribution for a class ($D_{s}$) by plotting the normalized values of the Euclidean distances. Once the distribution of the training data is known, the data samples are divided into quantiles ($Q^{s}$). Each quantile ($q$) is then examined for the number of samples and the quantiles with the lowest number of samples are oversampled using SMOTE [40] if there are sufficient samples. SMOTE creates synthetic samples using the original data and different variants of it have been developed [41] since the initial method was proposed. The initial method of SMOTE is used in our proposed approach to address the potential lack of samples in a given quantile. Otherwise, no oversampling is applied due to insufficient samples. After the samples are divided into quantiles and optionally oversampled, each data sample is scored using either density or point scoring and rearranged accordingly ($R_{s}$). Finally, the rearranged training data ($T$) is ready to be passed to one of the learning methods for training and evaluation.
DDCL consists of two types of scoring methods: a sample density based method and a Euclidean distance based method. These are referred to as DDCL-Density and DDCL-Point respectively and are detailed in the following sections.
TABLE I: Datasets used for evaluating DDCL.
Dataset Data Instances Attributes Class Type
Breast Cancer Wisconsin (Diagnostic) 569 31 Binary
Cancer 457 9 Binary
Haberman’s Survival 306 3 Binary
Liver Disorder 345 6 Binary
Pima Indians Diabetes 768 8 Binary
New-Thyroid 215 5 Multi-class
Diabetes 130 [42] 86556 46 Multi-class